### v1

# Domain-Specific Translation with Open-Source Large Language Models: Resource-Oriented Analysis

Abstract
[…]
In this work, we compare the domain-specific translation performance of open-source autoregressive decoder-only large language models (LLMs) with task-oriented machine translation (MT) models. Our experiments focus on the medical domain and cover four language pairs with varied resource availability: English-to-French, English-to-Portuguese, English-to-Swahili, and Swahili-to-English. Despite recent advancements, LLMs exhibit a clear gap in specialized translation quality compared to multilingual encoder-decoder MT models such as NLLB-200. In three out of four language directions in our study, NLLB-200 3.3B outperforms all LLMs in the size range of 8B parameters in medical translation. While fine-tuning LLMs such as Mistral and Llama improves their performance at medical translation, these models still fall short compared to fine-tuned NLLB-200 3.3B models. Our findings highlight the ongoing need for specialized MT models to achieve higher-quality domain-specific translation, especially in medium-resource and low-resource settings. As larger LLMs outperform their 8B variants, this also encourages pre-training domain-specific medium-sized LMs to improve quality and efficiency in specialized translation tasks.
[…]
Domain-Specific Translation with
Open-Source Large Language Models:
Resource-Oriented Analysis
[…]
1 Introduction
In recent years, there have been several advancements related to employing large language models (LLMs) in diverse natural language processing (NLP) tasks, including machine translation (MT). Autoregressive decoder-only models such as BLOOM (Le Scao et al.,, 2022), ChatGPT (Brown et al.,, 2020), Mistral (Jiang et al.,, 2023), and Llama (Touvron et al.,, 2023) –to mention but a few– have been welcomed by NLP researchers and practitioners as a new paradigm, due to their outstanding in-context learning feature. As its name implies, in-context learning allows an LLM to make use of extra context at inference time to improve its output (Brown et al.,, 2020). When it comes to MT, researchers investigated incorporating all sorts of context, such as translation pairs similar to the new source text (Agrawal et al.,, 2023; Moslem et al., 2023a, ; Vilar et al.,, 2023), terminology (Moslem et al., 2023c, ), and dictionary words (Ghazvininejad et al.,, 2023). Usually prompting the model, i.e. writing instructions in a natural language, is enough without extra model fine-tuning. Nevertheless, several researchers investigated fine-tuning LLMs for MT tasks, for both sentence-level translation (Alves et al.,, 2023; Xu et al.,, 2023), and adaptive translation (Moslem et al., 2023b, ; Zhang et al.,, 2023; Stap et al.,, 2024).
Despite these advances, there are still several challenges for using LLMs for MT, including:
  • clear gaps when it comes to language support, whether for the generic domain (Robinson et al.,, 2023), or for specialized domains (Moslem et al., 2023a, ).
  • challenges of deploying open-source LLMs in production given their massive sizes. The size of Llama 8B is more than double the size of NLLB-200 3.3B, while Llama 70B and Llama 405B are more than 20 times and 122 times larger, respectively.
This reopens the door to the question whether we still need conventional encoder-decoder MT models. In this work, we focus on domain-specific translation, and use NLLB-200 3.3B, both the baseline and fine-tuned models, as our main task-oriented benchmarks. The majority of decoder-only LLMs we explore are in the size range of 8B parameters, namely Mistral 7B (Jiang et al.,, 2023), Llama-3 8B, Llama-3.1 8B (Dubey et al.,, 2024), and Gemma 7B (Mesnard et al.,, 2024). Moreover, we evaluate Llama-3 70B, and Mixtral 8x7B, which is a mixture of 7B experts model, as well as the massive Llama-3.1 405B.
Even when comparing two languages with relatively higher data resources, such as French and Portuguese, we observe a considerable difference in the translation performance of these models. Moreover, for specialized translation, as in the case of the medical domain, many languages suffer from resource scarcity. This is clearly demonstrated by the quality of the English-to-Swahili and Swahili-to-English medical translation. Such resource differences between French, Portuguese, and Swahili are clearly reflected in their support levels by LLMs.
The main takeaways from our study include:
  • Domain-specific translation, e.g. in the medical domain, is challenging for general-purpose open-source LLMs, even for higher-resource languages.
  • In only one out of the four language directions included in our study, i.e. English-to-French translation, decoder-only LLMs in the size range of 8B parameters outperforms the encoder-decoder task-oriented NLLB-200 3.3B in zero-shot translation. For the other three language directions, English-to-Portuguese, English-to-Swahili, and Swahili-to-English, NLLB-200 3.3B excels at sentence-level translation.
  • The massive Llama-3.1 405B outperforms all the baseline models in the study. However, deploying Llama-3.1 405B in production can be challenging and inefficient. [^1] Quantized versions of Llama-3.1 405B require 3-4 H100 GPUs for one active user. Scaling this number to simultaneously serve several users can be too expensive for most medium-sized businesses. Moreover, latency can be counterproductive in translation settings. Hence, it can be used for offline data augmentation and knowledge distillation to smaller models.
  • Fine-tuning task-oriented MT models such as NLLB-200 3.3B can be effective when a medium-sized, domain-specific dataset is available. In such cases, using decoder-only LLMs with around 8B parameters may lead to diminishing returns.
The next sections elaborate on our methodology and evaluation of various open-source LLMs for domain-specific, sentence-level zero-shot translation and retrieval-augmented one-shot translation, as well as fine-tuning LLMs for both use cases.
2 Training Data
For our experiments, we used diverse publicly available datasets. This section elaborates on the data statistics and data preparation approaches.
2.1 French and Portuguese
We gathered medical data for English-to-Portuguese and English-to-French translations from the OPUS repository, specifically the ELRC, EMEA (Tiedemann,, 2012), SciELO (Soares et al.,, 2018), and TICO-19 (Anastasopoulos et al.,, 2020) datasets. Following a combination of rule-based and semantic filtering techniques, we curated 2,905,418 segments for English-to-Portuguese and 303,948 segments for English-to-French. Table 1 and Table 2 provide detailed breakdown of segment counts before and after filtering. [^2] Data preparation scripts: https://github.com/ymoslem/MT-Preparation/
Filtering
Dataset Raw Rule-Based Semantic
ELRC_2922 4,365 4,241 4,241
ELRC_2923 502 486 486
ELRC_3382 6,593 6,547 6,547
ELRC-antibiotic 1,072 1,051 1,051
ELRC-EUROPARL_covid 789 773 773
ELRC-presscorner_covid 11,361 11,296 11,296
ELRC-wikipedia_health 4,366 4,269 4,269
EMEA 1,092,568 313,240 313,240
Tico-19 3,071 2,971 2,971
Total 1,156,317 376,216 376,216
Global Filtering 303,948
Table 1: Medical datasets for French
Filtering
Dataset Raw Rule-Based Semantic
ELRC_2922 7,292 7,197 7,197
ELRC_2923 560 545 545
ELRC_3382 3,513 3,479 3,479
ELRC-antibiotic 839 821 821
ELRC-EUROPARL_covid 864 843 843
ELRC-presscorner_covid 6,319 6,263 6,263
ELRC-wikipedia_health 7,293 7,223 7,223
EMEA 1,082,144 310,619 310,619
SciELO 3,084,830 2,732,045 2,732,045
Tico-19 3,071 3,069 3,069
Total 4,196,725 3,072,104 3,072,104
Global Filtering 2,905,418
Table 2: Medical datasets for Portuguese
For the training process, we combined data from the aforementioned medical datasets and performed random sampling from the resulting large dataset. For each language pair, we selected 10,000 segments for a small training set, 100,000 segments for a medium-sized training set, and 10,000 segments for the test set in the medical domain. To retrieve fuzzy matches, we extracted semantically similar translation pairs from a context dataset. We sampled unique 50,000 segments for the context dataset of the small training dataset, and more 50,000 segments, making a context dataset with 100,000 for the medium training dataset.
2.2 Swahili
We collected medical and generic datasets for Swahili-to-English and English-to-Swahili models from OPUS (Tiedemann,, 2012). We used Global Voices and Wikimedia for the generic domain, and TICO-19 (Anastasopoulos et al.,, 2020) and Wikipedia articles for the medical domain. After applying rule-based and semantic filtering, we obtained 39,415 segments for the generic domain and 3,514 segments for the medical domain. The breakdown of segment numbers in our datasets before and after filtering is shown in Table 3.
In the generic domain, we used 10,000 segments for the test and 10,000 segments for the test context. Since the data we have for the medical domain is limited, we used only 1,000 segments for the test and 1,000 segments for the test context. For training, we mixed datasets from both the medical and generic domains. In total, we prepared 10,000 segments, with 9,000 from the generic domain and 1,000 from the medical domain. For the training context, we used all the remaining data, which includes 514 segments from the medical domain and 10,415 segments from the generic domain.
Filtering
Domain Dataset Raw Rule-Based Semantic
Medical ELRC-wikipedia_health 608 594 593
Tico-19 3,071 3,058 3,056
Total 3,679 3,652 3,649
Global Filtering 3,514
Generic GlobalVoices 32,307 29,633 29,623
Wikimedia 16,289 13,834 13,829
Total 48,596 43,467 43,452
Global Filtering 39,409
Table 3: Medical and generic datasets for Swahili
3 Experiments
In our experiments, we start by evaluating the baselines of various open-source models, including both autoregressive decoder-only LLMs, and the encoder-decoder NLLB-200 3.3 model, for both zero-shot and one-shot translation. Then, we fine-tune the models that show the best performance. In most cases, the LLMs chosen for fine-tuning are Mistral 7B and Llama-3 8B for their translation quality and efficiency. Following Moslem et al., 2023a , [^3] Code for the experiments: https://github.com/ymoslem/Adaptive-MT-LLM-Fine-tuning we fine-tune these LLMs using a combination of zero-shot and one-shot prompts in the medical domain. For the one-shot prompt experiments, we augment each translation pair with an additional translation example that is a fuzzy match to the source sentence. We extract fuzzy matches from a unique dataset which we refer to as the “context dataset” (cf. Section 2). Fuzzy matches are identified through calculating the semantic similarity between the embeddings of the source segments in the training datasets and those in the “context dataset”. The top match is selected for a one-shot prompt. Both the baseline and the fine-tuned models are then evaluated separately on zero-shot and one-shot prompts.
Fuzzy match retrieval involves three steps: (i) creating the embeddings, (ii) indexing, and (iii) retrieval. For creating the embeddings, we use Sentence-Transformers (Reimers and Gurevych,, 2019) with Microsoft Multilingual-MiniLM-L12-H384. Additionally, we tested other embedding models, and reported the results in Appendix A. For indexing, we use Faiss (Johnson et al.,, 2019), a library for efficient similarity search and clustering of dense vectors, training an IndexIVFFlat index with IndexFlatL2 as a quantizer. Retrieval was done by similarity search with cosine similarity.
In our experiments, we fine-tune a range of autoregressive decoder-only models, namely: Mistral 7B, Mixtral 7x8B, Llama-3 8B, and Gemma 7B. For efficiency, we employ QLoRA, a parameter-efficient fine-tuning technique. We fine-tune the models for one epoch, as in Moslem et al., 2023a , following the same configuration settings. As Llama-3 8B revealed the best results (cf. Section 4), we experimented with fine-tuning it with two sizes of data sets, (i) small [s+1], which is used for fine-tuning all the other models, and (ii) and medium [m+1]. The details of these datasets can be referred to in Section 2.
Moreover, we evaluated the NLLB-200 3.3B model, which is a state-of-the-art sequence-to-sequence machine translation model, based on the encoder-decoder Transformer architecture. For NLLB-200 3.3B fine-tuning, we conducted 4 experiments: (i) sentence-level fine-tuning with the small dataset [s]; (ii) sentence-level fine-tuning with the medium dataset [m]; (iii) concatenation of a fuzzy match with the new sentence pair in the small dataset [s+1]; and (iv) concatenation of a fuzzy match with the new sentence pair in the medium dataset [m+1]. The concatenated fuzzy segment and new segment are separated by both the language code and a special separator (Bulte and Tezcan,, 2019; Moslem et al., 2023b, ). At the decoding time, for the two models that use fuzzy-match concatenation, we apply teacher forcing of the fuzzy target as a translation prefix so that the model completes the translation of the new source accordingly. [^4] Refer to the CTranslate2’s “target_prefix” option.
For faster inference, we use vLLM for translation with all autoregressive models, while we use CTranslate2 for NLLB-200 3.3, which is not supported by vLLM. [^5] CTranslate2 supports some, but not all the autoregressive models in our experiments; hence, we decided to use vLLM for these models. Appendix A compares scores obtained from CTranslate2 and vLLM. Due to the massive size of Llama 3.1 405B, we had to use Activation-aware Weight Quantization (AWQ) (Lin et al.,, 2023), while other baseline models do not use quantization at inference time.
We conducted our experiments with both higher-resource (French and Portuguese) and lower-resource (Swahili) languages. French and Portuguese have extensive data available, including medical datasets. However, for Swahili, due to limited data, we created a small dataset in the medical domain and a generic dataset that includes data outside the medical domain. These datasets were partitioned into training and test sets, with the remaining data designated as the context dataset (cf. Section 2).
Moreover, we ran several additional experiments to find the best set up, including embedding models, learning rates, and inference engines. Appendix A briefly summarizes these experiments.
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
NLLB-200 3.3B baseline zero-shot 41.37 64.10 52.11 59.32
one-shot 43.04 65.55 54.49 55.84
finetuned [s] zero-shot 44.22 66.94 50.09 75.40
finetuned [m] zero-shot 49.56 70.96 44.97 81.36
finetuned [s+1] zero-shot 45.20 68.25 48.38 77.07
one-shot 47.93 69.15 46.12 77.95
finetuned [m+1] zero-shot 53.13 73.49 41.61 83.83
one-shot 53.67 72.74 42.28 81.19
Gemma 7B baseline zero-shot 26.21 48.31 73.19 33.91
one-shot 47.97 67.78 47.35 71.64
finetuned [s+1] zero-shot 41.08 63.74 54.59 68.79
one-shot 41.42 65.87 51.08 48.94
Llama-3 8B baseline zero-shot 34.40 58.08 59.82 59.36
one-shot 48.45 67.79 46.64 72.80
finetuned [s+1] zero-shot 43.71 64.93 50.62 73.05
one-shot 51.11 69.53 43.82 77.88
finetuned [m+1] zero-shot 49.67 69.15 44.97 77.87
one-shot 54.77 72.00 40.56 79.75
Llama-3.1 8B baseline zero-shot 33.89 57.67 59.49 58.58
one-shot 48.58 67.96 46.30 73.13
Llama-3 70B baseline zero-shot 39.74 62.51 53.36 66.75
one-shot 51.93 70.57 43.03 76.76
Llama-3.1 405B baseline zero-shot 42.61 64.92 50.48 69.56
one-shot 53.61 71.83 41.29 79.32
Mistral 7B baseline zero-shot 30.61 54.79 62.14 48.56
one-shot 46.18 66.12 48.02 69.13
finetuned [s+1] zero-shot 39.46 64.63 50.98 71.84
one-shot 48.40 69.20 44.57 76.64
BioMistral baseline zero-shot 24.13 51.39 68.68 42.10
one-shot 43.41 63.71 51.79 63.81
Mixtral 8x7B baseline zero-shot 38.06 61.17 55.27 62.29
one-shot 50.98 69.79 43.74 75.57
finetuned [s+1] zero-shot 44.32 66.19 48.28 75.50
one-shot 47.29 69.10 43.95 78.74
Table 4: Evaluation of LLMs for English-to-French medical translation. After fine-tuning Llama-3 8B and NLLB-200 3.3 on the medium-sized dataset with fuzzy matches, they outperform all the evaluated models (bold). Llama-3.1 405B (one-shot) outperforms all the “baseline” models, while Llama-3 8B and Llama-3.1 8B (one-shot) are the best among baseline LLMs in the size range of 8B parameters. Nevertheless, for zero-shot translation, NLLB-200 3.3B is still the top-performer among all the evaluated baseline pre-trained models, except Llama-3.1 405B.
4 Evaluation and Discussion
In this section, we discuss the evaluation results of our experiments. We experimented with both pure prompting LLMs for the task of translation, and fine-tuning these baselines first with zero-shot and one-shot translation prompts. To evaluate our systems, we calculated BLEU (Papineni et al.,, 2002), chrF++ (Popović,, 2017), and TER (Snover et al.,, 2006), as implemented in the sacreBLEU library (Post,, 2018). [^6] https://github.com/mjpost/sacrebleu For semantic evaluation, we used COMET Rei et al., (2020). [^7] https://github.com/Unbabel/COMET, [^8] In particular, we used the “wmt20-comet-da” model. As COMET does not fully support Swahili, we used AfriCOMET-MTL (multitask learning) (Wang et al.,, 2024) evaluation model. [^9] https://huggingface.co/masakhane/africomet-mtl
In this section, we explore the medical translation performance of a range of open-source models, including, Mistral, Mixtral (Jiang et al.,, 2023), Llama 3/3.1 (Dubey et al.,, 2024), and Gemma, for both zero-shot and one-shot translation. For each language, we evaluate baselines (without fine-tuning). For models that show better performance, we also experimented with fine-tuning these LLMs, and hence we report the evaluation results of the fine-tuned versions too. Moreover, we evaluated NLLB-200 3.3 baseline and four fine-tuned variants (cf. Section 3).
4.1 French and Portuguese
In this section, we elaborate on the evaluation results for English-to-French and English-to-Portuguese medical translation (cf. Table 4 and Table 5) for a range of open-source models.
4.1.1 Baseline Evaluation
For both French and Portuguese, one-shot translation, i.e. prepending the prompt with a similar translation pair at inference time, outperforms zero-shot translation, using the baselines of all the covered autoregressive decoder-only LLMs.
For zero-shot translation, NLLB-200 3.3B, the encoder-decoder MT model, outperforms all the baseline models in the size range of 8B parameters, while Llama-3 8B and Llama-3.1 one-shot translation outperforms the models in the same size range. Although Llama-3.1 models are promoted as being “multilingual” (Dubey et al.,, 2024), Llama-3 8B and Llama 3.1 8B show comparable performance.
When it comes to larger models, Mixtral 8x7B outperforms Llama 3 8B. Nevertheless, the gigantic Llama-3 70B and Llama-3.1 405B achieve the best performance among all the baseline models, especially with one-shot translation.
BioMistral (Labrak et al.,, 2024) is an open-source LLM tailored for the biomedical domain, utilizing Mistral as its foundation model and further pre-trained on PubMed Central. It supports a range of languages, including English, French, German, Dutch, Spanish, Portuguese, Polish, and Romanian. However, in our experiments for the medical domain, Mistral outperforms BioMistral at both zero-shot and one-shot English-to-French translation.
Finally, Gemma 7B has the worst performance among all the models we evaluated for zero-shot translation. However, for one-shot translation, Gemma slightly outperforms Mistral.
4.1.2 Fine-tuning Evaluation
We fine-tuned Mistral 7B, Mixtral 7x8B, Llama-3 8B, and Gemma 7B. In this section, we elaborate on our evaluation results.
Decoder-only LLMs:
After fine-tuning Mistral, we observe considerable performance gains in both zero-shot and one-shot translation for both French and Portuguese (cf. Table 4 and Table 5). Interestingly, while fine-tuning Mixtral 8x7B enhancs its evaluation scores for zero-shot translation, it degrades its one-shot translation performance. Fine-tuning both Mistral and Mixtral 8x7B, zero-shot translation of Mixtral 8x7B still outperforms Mistral 7B, while we do not observe the same level of improvement for one-shot translation. Fine-tuning Gemma 7B considerably improves zero-shot translation, while degrades one-shot English-to-French translation. Nevertheless, except for BLEU, Mistral still outperforms Gemma after fine-tuning both models for both zero-shot and one-shot translation based on all other metrics. The fine-tuned Llama-3 8B for English-to-French medical translation outperforms Llama-3 70B baseline (i.e. without fine-tuning) for zero-shot translation, while the evaluation scores for one-shot translation are relatively comparable. For English-to-Portuguese medical translation, the larger model Llama-3 70B still outperforms the fine-tuned Llama-3 8B model.
NLLB-200 3.3:
Fine-tuning the task-oriented encoder-decoder MT model, NLLB-200 3.3, reveals interesting findings. After fine-tuning on the medical dataset, NLLB-200 3.3 outperforms the fine-tuned Llama-3 8B at all the experiments in Portuguese (cf. Table 4), while showing comparable results in French (Table 5). For French, NLLB-200 3.3 fine-tuned on the medium-sized dataset with concatenated fuzzy matches [m+1] (cf. Section 2) outperforms the gigantic Llama-3.1 405B. For Portuguese, the best results are achieved when NLLB-200 3.3 is fine-tuned with a regular sentence-level medium-sized dataset [m], which outperforms both the baseline and fine-tuned Llama-3 8B models.
It is worth mentioning that NLLB-200 3.3 was not originally pre-trained with fuzzy-match concatenation. Hence, one-shot translation might have quality issues in some translations, alleviated by fine-tuning. In other words, pre-training an MT model with domain-specific fuzzy-match concatenation can be considered (Bulte and Tezcan,, 2019) if this feature is required later for fine-tuning and inference. Moreover, using quality estimation tools such as MetricX (Juraska et al.,, 2024) is recommended to select the best translation from zero-shot (sentence-level) and one-shot (fuzzy-match concatenated) translations. Nevertheless, it is important to evaluate the quality gains of fuzzy-match concatenation per se. Fuzzy-match concatenation diversifies and increases the data, which might explain improvements in zero-shot translation, but not one-shot translation. As in the case of Portuguese and Swahili (cf. Section 4.2), conventional sentence-level fine-tuning can be sufficient. Other data augmentation techniques can be considered, including monolingual data back-translation (Sennrich et al.,, 2016; Edunov et al.,, 2018) and domain-specific text generation (Moslem et al.,, 2022).
In a nutshell, when a small-to-medium-sized domain-specific dataset is available, fine-tuning the task-oriented NLLB-200 3.3 achieves comparable or better results than fine-tuning LLMs in the size range of 8B parameters such as Llama-3 8B and Mistral 7B, according to the ChrF++ and COMET scores. In the absence of a domain-specific dataset, one-shot translation with decoder-only models, especially larger LLMs such as Mixtral 8x7B, Llama-3 70B and Llama-3.1 405B, outperforms all the baseline pre-trained models.
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
NLLB-200 3.3B baseline zero-shot 39.32 64.98 48.19 81.77
one-shot 39.79 64.36 48.38 79.04
finetuned [s] zero-shot 43.86 67.52 43.43 85.14
finetuned [m] zero-shot 46.41 69.13 41.50 87.29
finetuned [s+1] zero-shot 44.22 67.82 43.19 85.71
one-shot 43.74 67.48 43.58 85.34
finetuned [m+1] zero-shot 46.29 69.08 41.44 87.36
one-shot 45.90 68.89 41.66 87.15
Gemma 7B baseline zero-shot 33.45 57.91 58.69 58.13
one-shot 40.39 65.02 47.41 80.59
finetuned [s+1] zero-shot 39.72 64.10 47.88 79.11
one-shot 40.28 64.57 47.09 79.98
Llama-3 8B baseline zero-shot 38.68 63.39 49.28 77.68
one-shot 41.15 65.49 46.24 82.27
finetuned [s+1] zero-shot 42.10 66.01 45.20 82.48
one-shot 42.27 66.07 45.26 82.53
finetuned [m+1] zero-shot 43.63 67.15 43.89 84.22
one-shot 44.00 67.37 43.43 84.54
Llama-3.1 8B baseline zero-shot 38.77 63.51 49.16 78.17
one-shot 41.44 65.65 46.14 82.47
Llama-3 70B baseline zero-shot 43.86 67.35 43.95 84.05
one-shot 45.95 68.82 41.68 86.58
Llama-3.1 405B baseline zero-shot 46.40 69.08 41.92 86.05
one-shot 47.76 70.01 40.15 88.07
Mistral 7B baseline zero-shot 33.60 60.08 55.23 63.52
one-shot 37.64 62.92 49.35 78.05
finetuned [s+1] zero-shot 40.24 64.77 46.96 80.23
one-shot 40.62 64.97 46.84 80.98
Mixtral 8x7B baseline zero-shot 39.58 64.06 48.07 77.46
one-shot 42.31 66.33 45.12 83.59
Table 5: Evaluation of LLMs for English-to-Portuguese medical translation. Llama-3.1 405B (one-shot) outperforms all the models (bold), while Llama-3 8B and Llama-3.1 8B (one-shot) are the best among LLMs in the size range of 8B parameters. After fine-tuning Llama-3 8B, it outperforms the decoder-only baseline models in the size range of 8B parameters a well as the fine-tuned Mistral, while Llama-3 70B and Llama-3.1 405B still outperform the fine-tuned Llama-3 8B. Nevertheless, for zero-shot translation, NLLB-200 3.3B is still the top-performer among all the evaluated pre-trained models in the size range of 8B parameters. After fine-tuning NLLB-200 3.3B with both the medium dataset (finetuned [m]), it outperforms all the models in the size range of 8B parameters (underlined). Fine-tuning NLLB-200 3.3B on the sentence-level medium dataset (finetuned [m]) outperform fine-tuning it on the medium dataset concatenated with fuzzy matches (finetuned [m+1]).
4.2 Swahili
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ AfriCOMET ↑
NLLB-200 3.3B baseline zero-shot 44.58 64.73 42.71 84.11
one-shot 43.66 64.03 42.85 83.92
finetuned zero-shot 45.80 65.75 40.95 83.98
finetuned [+1] zero-shot 47.11 66.70 40.23 84.29
one-shot 45.64 65.85 41.18 83.92
Gemma 7B baseline zero-shot 35.20 58.71 55.80 81.70
one-shot 38.98 61.37 49.78 83.02
Llama-3 8B baseline zero-shot 29.19 53.35 68.78 79.77
one-shot 35.91 57.86 53.75 81.60
finetuned [+1] zero-shot 38.33 60.11 51.09 82.16
one-shot 39.47 60.83 49.83 82.24
Llama-3.1 8B baseline zero-shot 28.19 51.61 66.28 78.06
one-shot 33.92 56.11 56.14 80.54
Llama-3 70B baseline zero-shot 38.31 61.14 56.23 82.27
one-shot 44.06 64.31 45.83 83.36
Llama-3.1 405B baseline zero-shot 47.76 66.57 44.06 83.12
one-shot 52.00 68.79 39.77 83.23
Mistral 7B baseline zero-shot 11.94 32.67 97.07 59.12
one-shot 16.74 36.83 85.52 65.26
finetuned [+1] zero-shot 31.62 52.79 58.44 77.67
one-shot 32.50 53.73 56.97 78.59
Mixtral 8x7B baseline zero-shot 19.32 41.50 82.05 69.85
one-shot 24.09 45.04 70.75 73.23
Table 6: Evaluation of LLMs for Swahili-to-English medical translation. Llama-3.1 405B (one-shot) is the top-performing model (bold). Nevertheless, the encoder-decoder MT model, NLLB-200 3.3B (both baseline and fine-tuned) is the top-performer among the models in the size range of 8B parameters (underlined).
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ AfriCOMET ↑
NLLB-200 3.3B baseline zero-shot 42.87 61.59 48.69 82.00
one-shot 41.38 59.99 50.85 80.92
finetuned zero-shot 43.67 62.06 47.12 82.07
finetuned [+1] zero-shot 44.25 62.49 46.50 82.35
one-shot 43.31 61.72 47.18 82.04
Gemma 7B baseline zero-shot 32.70 54.87 66.20 79.00
one-shot 37.25 58.01 56.81 80.53
Llama-3 8B baseline zero-shot 29.24 52.97 76.90 78.28
one-shot 36.70 56.84 57.61 79.99
finetuned [+1] zero-shot 39.40 58.77 53.40 80.74
one-shot 40.18 59.04 52.14 80.93
Llama-3.1 8B baseline zero-shot 28.62 51.79 76.94 76.90
one-shot 35.09 55.49 60.30 78.92
Llama-3 70B baseline zero-shot 36.75 59.02 62.49 81.05
one-shot 42.27 61.52 51.12 81.84
Llama-3.1 405B baseline zero-shot 42.21 61.87 53.94 81.70
one-shot 45.61 63.46 47.84 82.10
Mistral 7B baseline zero-shot 17.04 36.69 93.06 59.83
one-shot 21.44 40.16 80.47 63.27
finetuned [+1] zero-shot 34.64 53.02 58.11 76.04
one-shot 35.08 53.36 57.28 76.24
Mixtral 8x7B baseline zero-shot 24.87 45.14 77.44 70.30
one-shot 28.53 47.98 68.75 72.46
Table 7: Evaluation of LLMs for Swahili-to-English generic translation. Llama-3.1 405B (one-shot) is the top-performing model (bold). Nevertheless, the encoder-decoder MT model, NLLB-200 3.3B (both baseline and fine-tuned) is the top-performer among the models in the size range of 8B parameters (underlined).
Based on the evaluation results for both the generic and medical domain, the top-performing model for Swahili-to-English translation continues to be NLLB-200 3.3, a Transformer-based encoder-decoder model specifically developed for MT tasks.
For both the medical domain (cf. Table 6) and the generic domain (cf. Table 7) of Swahili-to-English translation, Gemma and Llama-3 8B (especially after fine-tuning) are the top performers among autoregressive decoder-only models in the size range of 8B parameters. However, their performance still lags behind that of NLLB-200 3.3B. Nevertheless, Llama-3 70B evaluation scores are relatively comparable to those of NLLB-200 3.3B.
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ AfriCOMET ↑
NLLB-200 3.3B baseline zero-shot 32.61 58.24 53.29 82.85
one-shot 34.06 59.12 51.57 82.37
finetuned zero-shot 36.21 61.35 49.81 83.92
finetuned [+1] zero-shot 37.31 62.37 48.70 84.18
one-shot 37.52 62.29 48.38 83.93
Gemma 7B baseline zero-shot 11.66 29.87 108.69 60.62
one-shot 21.37 47.15 74.56 71.57
Llama-3 8B baseline zero-shot 10.10 33.82 116.26 60.95
one-shot 16.08 42.38 93.42 67.03
finetuned [+1] zero-shot 21.50 47.60 76.71 69.94
one-shot 23.58 48.99 71.09 71.69
Llama-3.1 8B baseline zero-shot 6.33 28.06 149.36 51.69
one-shot 11.64 36.91 112.11 58.98
Llama-3 70B baseline zero-shot 24.34 49.49 68.19 77.97
one-shot 28.06 53.67 59.98 79.87
Llama-3.1 405B baseline zero-shot 29.65 55.19 56.46 80.21
one-shot 31.65 57.27 54.95 81.41
Mistral 7B baseline zero-shot 0.64 11.02 309.37 25.79
one-shot 1.20 14.67 320.65 22.72
finetuned [+1] zero-shot 12.88 37.76 99.67 46.65
one-shot 14.83 39.81 91.12 49.91
Mixtral 8x7B baseline zero-shot 2.17 18.34 196.42 31.47
one-shot 4.00 24.88 184.27 36.66
Table 8: Evaluation of LLMs for the English-to-Swahili medical translation. NLLB-200 3.3B (both baseline and fine-tuned) is the top-performer among all the models, even those with larger sizes
Model Model Type Context BLEU ↑ ChrF++ ↑ TER ↓ AfriCOMET ↑
NLLB-200 3.3B baseline zero-shot 35.57 57.55 54.91 83.90
one-shot 33.79 55.93 56.34 83.23
finetuned zero-shot 38.98 60.21 52.00 85.05
finetuned [+1] zero-shot 39.58 60.78 51.39 85.45
one-shot 39.09 60.34 51.65 85.03
Gemma 7B baseline zero-shot 8.32 21.51 118.26 54.88
one-shot 22.51 45.37 81.91 71.08
Llama-3 8B baseline zero-shot 13.25 34.34 114.65 61.68
one-shot 20.03 43.80 89.42 69.56
finetuned [+1] zero-shot 26.47 49.24 74.29 73.74
one-shot 28.30 50.15 69.51 74.40
Llama-3.1 8B baseline zero-shot 10.17 30.44 138.65 55.64
one-shot 16.10 39.54 108.04 62.61
Llama-3 70B baseline zero-shot 25.77 48.20 73.21 77.67
one-shot 31.58 53.82 60.90 81.71
Llama-3.1 405B baseline zero-shot 32.53 53.98 59.74 80.82
one-shot 35.07 57.06 55.76 83.57
Mistral 7B baseline zero-shot 1.82 13.15 289.67 30.15
one-shot 2.70 17.02 320.44 27.88
finetuned [+1] zero-shot 19.28 42.12 96.24 53.35
one-shot 21.04 43.46 89.24 55.01
Mixtral 8x7B baseline zero-shot 5.35 20.89 157.16 40.93
one-shot 7.73 28.52 161.96 43.73
Table 9: Evaluation of LLMs for the English-to-Swahili generic translation. NLLB-200 3.3B (both baseline and fine-tuned) is the top-performer among all the models, even those with larger sizes
As expected, the evaluation results for Swahili-to-English translation with LLMs are superior to those of English-to-Swahili translation due to the limited support for lower-resource languages such as Swahili. Similar to Swahili-to-English translation, the top model for English-to-Swahili is still the NLLB-200 3.3. For both the medical domain (cf. Table 8) and the generic domain (cf. Table 9) of English-to-Swahili translation, all the autoregressive decoder-only models in our experiments demonstrate a lower translation quality. Even after fine-tuning Llama-3 8B, its evaluation scores are not on par with those of NLLB-200 3.3.
Although Llama 3.1 8B is promoted as a multilingual model Dubey et al., (2024), Llama 3 8B outperforms it for English-to-Swahili translation. Comparing model sizes, Llama 3 70B outperforms the smaller version of the model Llama 3 8B for both zero-shot and one-shot English-to-Swahili translation.
Fine-tuning:
We fine-tuned Mistral 7B and Llama-3 8B for both generic and medical Swahili-to-English and English-to-Swahili translation. While fine-tuning has improved the evaluation scores compared to the baselines, fine-tuning NLLB-200 3.3 outperforms even Llama-3 70B for Swahili-to-English translation and both Llama-3 70B and Llama-3.1 405B for English-to-Swahili translation.
5 Conclusions
After a few years of research on leveraging open-source LLMs for direct translation among other tasks, several businesses have been investigating the feasibility of integrating such general-purpose models into their specialized translation workflows. Among the main advantages of employing the in-context learning feature of LLMs is controlling consistency, domain terminology and style throughout the whole document or project Moslem et al., 2023a ; Wang et al., (2023). This can be achieved through feeding an LLM with similar translation pairs, pre-approved terms, or other useful data at inference time, hoping that the LLM’s output would be influenced by such provided linguistic information. Nevertheless, deploying open-source LLMs into production poses several efficiency challenges, including infrastructure requirements and latency at translation time. Moreover, LLMs show inconsistent quality across languages and domains. This raises the question of how conventional encoder-decoder MT models remain relevant, particularly for domain-specific translation.
As stated in Section 3, we evaluated a range of pre-trained models for medical translation. In addition, we fine-tuned selected models for zero-shot and one-shot translation. Our experiments covered both the multilingual encoder-decoder MT model, NLLB-200 3.3, and several autoregressive decoder-only LLMs, such as Gemma, Llama, Mistral, and Mixtral.
For French and Portuguese, some baselines of decoder-only LLMs, such as Llama-3 8B and Mixtral 8x7B, outperform NLLB-200 3.3. However, after fine-tuning NLLB-200 3.3, the performance gap narrows considerably, and NLLB becomes as good as or even surpasses the decoder-only LLMs, especially models in the size range of 8B parameters. For Swahili, only the massive Llama 3.1 405B model achieves performance comparable to NLLB-200 3.3, while all other evaluated decoder-only LLMs fall short.
Our results demonstrate the diminishing returns of using decoder-only LLMs in the size range of 8B parameters for domain-specific translation, especially when a small-to-medium-sized dataset is available for fine-tuning task-oriented translation models like NLLB-200 3.3. Larger LLMs, e.g. Llama-3 70B, Llama 3.1 405, or Mixtral 8x7B, can be considered for some use cases such as real-time adaptive translation or automatic post-editing, as they offer better quality than their 8B-parameter variants. However, their massive sizes make them less efficient for real-time scenarios in production, and more suitable for offline data augmentation use cases, or knowledge distillation into more efficient models. This presents a valuable opportunity for businesses to develop in-house medium-sized LMs, custom-built for their linguistic needs, especially for domain-specific and low-resource scenarios. In the meantime, task-oriented encoder-decoder MT models remain a core component in high-quality domain-specific translation workflows.
Appendix A Appendix
Lang Model Domain Context Embedding Model BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-PT baseline Medical zero-shot N/A 32.70 59.81 56.39 59.92
one-shot all-MiniLM-L6-v2 37.43 62.89 49.44 78.01
all-mpnet-base-v2 37.49 62.82 49.33 78.25
Multilingual-MiniLM-L12-H384 37.41 62.81 49.47 77.91
finetuned Medical zero-shot all-MiniLM-L6-v2 39.89 64.39 47.16 79.77
all-mpnet-base-v2 39.44 64.05 47.95 79.22
Multilingual-MiniLM-L12-H384 40.05 64.68 47.08 80.12
one-shot all-MiniLM-L6-v2 40.58 64.92 46.49 81.18
all-mpnet-base-v2 40.20 64.57 47.32 80.59
Multilingual-MiniLM-L12-H384 40.39 64.86 47.11 80.61
EN-SW baseline Medical zero-shot N/A 0.49 10.22 290.16 25.34
one-shot all-MiniLM-L6-v2 1.48 14.67 261.35 24.09
Multilingual-MiniLM-L12-H384 0.97 13.18 283.04 21.87
Generic zero-shot N/A 1.30 11.50 278.97 28.16
one-shot all-MiniLM-L6-v2 2.26 15.17 278.88 27.15
Multilingual-MiniLM-L12-H384 1.99 14.45 294.13 25.65
finetuned Medical zero-shot all-MiniLM-L6-v2 8.83 33.11 140.96 42.44
Multilingual-MiniLM-L12-H384 13.38 38.06 96.06 46.49
one-shot all-MiniLM-L6-v2 11.10 36.81 128.07 47.81
Multilingual-MiniLM-L12-H384 15.67 40.67 84.47 49.57
Generic zero-shot all-MiniLM-L6-v2 12.21 34.95 147.31 47.47
Multilingual-MiniLM-L12-H384 19.12 41.71 94.07 52.85
one-shot all-MiniLM-L6-v2 13.97 37.32 133.67 50.51
Multilingual-MiniLM-L12-H384 21.12 43.16 85.21 54.37
Table 10: Comparison of embedding models for English-to-Portuguese (EN-PT) and English-to-Swahili (EN-SW) translation, showing baseline and fine-tuned performance on medical and generic test datasets.
In this appendix, we briefly summarize the results of extra experiments we conducted. The experiments include various embedding models used to create fuzzy matches, the impact of learning rate choice, dataset packing, and comparing inference engines, CTranslate2 and vLLM. While the results in the main paper do not use quantization, some results in the appendix use quantization for faster experimental turnaround, which is clarified in each section. Moreover, we compare quantization approaches in Section A.5. We use the same evaluation metics BLEU, ChrF++, TER, and COMET. For Swahili, the COMET column shows AfriCOMET-MTL scores.
A.1 Embedding Models
To create embeddings used later to generate the retrieval index (cf. Section 3), we mainly used MicrosoftMultilingual-MiniLM-L12-H384” throughout our paper, as it is a multilingual model. Furthermore, we experimented with “all-MiniLM-L6-v2” and “all-mpnet-base-v2”, which are an efficient embedding models for English. As we retrieve translation pairs based on the source text, we experimented with English-to-Portuguese and English-to-Swahili language directions. Table 10 shows evaluation of the Mistral 7B model, both the baseline and fine-tuned version, using CTranslate2 with int8 quantization. For English-to-Portuguese, the results were comparable, while for English-to-Swahili the results achieved by using Microsoft “Multilingual-MiniLM-L12-H384” were generally better than the two other embedding models, so we used it for all our experiments.
A.2 Learning Rate
Before conducting the main experiments, we tested a range of learning rate values to identify the optimal value. Translations were generated using CTranslate2 without quantization.
A.2.1 Mistral
The learning rates we explored while fine-tuning Mistral 7B include 3e-4, 1e-3 and 2e-3. The results are presented in Table 11. Overall, the learning rate of 1e-3 produced the highest performance. For the sake of consistency, we then used the learning rate 1e-3 for fine-tuning all the decoder-only LLMs.
A.2.2 NLLB
Similarly, when fine-tuning NLLB-200 3.3, we tried a few learning rates. Unlike Mistral, the learning rate 1e-3 did not achieve the best result. However, the learning rate 5e-5 revealed better evaluation scores, as illustrated by Table 12. Hence, we used it when fine-tuning NLLB-200 3.3 for all the language pairs.
Lang Domain Context LR BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot 1e-3 39.33 64.57 51.18 71.89
2e-3 41.73 64.45 50.15 71.47
one-shot 1e-3 45.31 68.57 45.81 74.04
2e-3 48.15 68.94 44.55 76.21
EN-PT medical zero-shot 1e-3 40.05 64.68 47.08 80.12
2e-3 38.63 63.33 49.55 77.30
1e-4 38.90 63.80 48.08 78.90
3e-4 39.82 64.46 47.14 80.06
one-shot 1e-3 40.39 64.86 47.11 80.61
2e-3 39.00 63.63 48.74 78.84
1e-4 39.43 64.11 47.70 79.89
3e-4 40.30 64.78 46.86 80.91
EN-SW medical zero-shot 1e-3 13.38 38.06 96.06 46.49
2e-3 11.61 36.35 113.01 45.55
3e-4 4.96 25.29 162.21 31.13
one-shot 1e-3 15.67 40.67 84.47 49.57
2e-3 14.46 39.46 96.04 48.70
3e-4 6.65 28.24 147.95 35.71
generic zero-shot 1e-3 19.12 41.71 94.07 52.85
2e-3 19.70 41.55 91.85 53.63
3e-4 11.21 32.51 136.13 40.66
one-shot 1e-3 21.12 43.16 85.21 54.37
2e-3 21.66 42.88 83.99 54.98
3e-4 12.53 34.17 129.39 42.99
SW-EN medical zero-shot 1e-3 31.81 52.87 58.53 77.38
2e-3 27.03 47.81 63.38 73.69
one-shot 1e-3 32.25 53.50 57.52 78.46
2e-3 27.44 48.21 63.37 74.89
generic zero-shot 1e-3 34.49 52.93 58.24 75.87
2e-3 31.97 50.17 61.07 73.84
one-shot 1e-3 34.97 53.28 57.21 76.06
2e-3 32.29 50.45 60.71 74.22
Table 11: Comparison of learning rates (LR) when fine-tuning Mistral-7B model across language pairs, showing zero-shot and one-shot performance.
Lang Domain LR BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical 1e-3 26.70 51.82 74.46 27.97
1e-4 43.40 66.46 50.99 74.40
2e-4 41.53 65.25 53.01 71.01
4e-5 42.27 66.70 50.40 75.01
5e-5 44.02 67.08 50.24 75.34
6e-5 42.85 66.86 50.33 75.16
7e-5 43.76 66.77 50.69 74.40
EN-SW medical 1e-3 20.94 46.24 72.65 66.03
1e-4 36.32 61.57 49.38 83.83
5e-5 36.30 61.31 49.48 83.99
generic 1e-3 21.88 45.49 78.23 67.72
1e-4 38.08 59.57 52.53 84.70
5e-5 38.90 60.14 51.98 85.08
Table 12: Comparison of learning rates (LR) when fine-tuning NLLB-3.3 model for zero-shot (sentence-level) translation.
A.3 Dataset Packing
This section examines the performances of the models with and without packing. Packing dataset, initially introduced by (Raffel et al.,, 2020), helps increase the training efficiency by enabling fine-tuning on more sequences. In this method, multiple short examples are packed in the same input sequence. Packing tokens into the maximum sequence length helps improve training efficiency. While packing is primarily intended to improve the training throughput, our experiments demonstrate it also improves model performance. In our experiments, we use the HuggingFace’s SFTTrainer, which supports example packing.
In all the experiments in the main paper, we use packing=True for the SFTTrainer. Hence, we set add_eos_token to False as the trainer’s dataset packing is True. By default, SFTTrainer adds eos_token if dataset packing=True, so there is no need to add it manually. If packing is False, then add_eos_token should be set to True. The evaluation results use CTranslate2 with int8 quantization and are summarized in Table 13.
Lang Domain Context Packing BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot False 34.75 61.45 65.37 67.74
True 39.33 64.57 51.18 71.89
one-shot False 17.53 53.53 195.57 37.93
True 45.31 68.57 45.81 74.04
EN-PT medical zero-shot False 39.39 64.32 47.30 79.15
True 40.05 64.68 47.08 80.12
one-shot False 39.70 64.49 47.26 79.63
True 40.39 64.86 47.11 80.61
EN-SW medical zero-shot False 18.04 44.15 81.44 56.30
True 13.38 38.06 96.06 46.49
one-shot False 12.27 42.26 121.31 54.45
True 15.67 40.67 84.47 49.57
generic zero-shot False 24.25 46.07 75.28 60.51
True 19.12 41.71 94.07 52.85
one-shot False 10.30 39.53 172.85 54.83
True 21.12 43.16 85.21 54.37
SW-EN medical zero-shot False 30.12 52.96 63.28 77.12
True 31.81 52.87 58.53 77.38
one-shot False 9.81 41.21 252.60 67.72
True 32.25 53.50 57.52 78.46
generic zero-shot False 34.42 53.49 60.02 76.38
True 34.49 52.93 58.24 75.87
one-shot False 6.15 31.94 459.91 60.20
True 34.97 53.28 57.21 76.06
Table 13: Comparison of packing settings when fine-tuning Mistral.
A.4 CTranslate2 vs vLLM
CTranslate2 and vLLM are both efficient inference engines for LLMs. While they have an overlapping list of supported models, some models are not supported by both of them. For example, NLLB-200 models are supported by CTranslate2, but not supported by vLLM. At the time of writing this paper, Mixtral 8x7B is not supported by CTranslate2. For the sake of consistency of our results, all the models reported in the main paper use vLLM, except NLLB-200 3.3. Nevertheless, in this section, we show a brief comparison between CTranslate2 and vLLM evaluation results. The evaluation results are generally comparable between the two inference engines when evaluating Mistral and NLLB.
Lang Domain Context Library BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot CTranslate2 30.54 54.76 62.27 48.89
vLLM 30.61 54.79 62.14 48.56
one-shot CTranslate2 46.18 66.18 47.59 69.25
vLLM 46.18 66.12 48.02 69.13
EN-PT medical zero-shot CTranslate2 33.53 60.09 55.39 63.22
vLLM 33.60 60.08 55.23 63.52
one-shot CTranslate2 37.63 62.92 49.26 77.99
vLLM 37.64 62.92 49.35 78.05
EN-SW medical zero-shot CTranslate2 0.67 10.88 312.19 25.89
vLLM 0.64 11.02 309.37 25.79
one-shot CTranslate2 1.09 13.85 326.60 22.35
vLLM 1.20 14.67 320.65 22.72
generic zero-shot CTranslate2 1.72 12.74 289.67 29.64
vLLM 1.82 13.15 289.67 30.15
one-shot CTranslate2 2.49 15.96 320.44 27.19
vLLM 2.70 17.02 320.44 27.88
SW-EN medical zero-shot CTranslate2 11.84 33.31 98.46 60.11
vLLM 11.94 32.67 97.07 59.12
one-shot CTranslate2 16.86 37.97 82.50 66.41
vLLM 16.74 36.83 85.52 65.26
generic zero-shot CTranslate2 16.77 36.99 95.29 60.03
vLLM 17.04 36.69 93.06 59.83
one-shot CTranslate2 21.83 40.62 79.75 63.53
vLLM 21.44 40.16 80.47 63.27
Table 14: Evaluation of Mistral 7B performance (baseline model without quantization) using CTranslate2 and vLLM inference engines, showing comparable results.
A.5 Quantization
We compared three cases of quantization when working with NLLB-200 CTranslate2, i.e. int8, float16, and no quantization (cf. Table 15). Models with int8 quantization have the smallest size, and they excel on CPU setups, while models with float16 quantization are smaller than non-quantized models, but larger than int8 models. Float16 models are faster on GPU setups. When it comes to translation performance, the differences between the three setups are marginal. All NLLB-200 3.3 results in the main paper are for non-quantized models.
Moreover, we compared generation with Llama-3 70B with and without Activation-aware Weight Quantization (AWQ) using vLLM. As Table 16 shows, results without quantization are slightly better. All the results in the main paper for LLMs are for non-quantized models, except Llama-3.1 405B, due to commutating constrains. Hence, we tested Llama-3 70B with and without quantization and reported the results in this section to understand the effect of AWQ on the translation quality.
Lang Domain Context Quantization BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot None 41.37 64.10 52.11 59.32
Float16 41.30 64.08 52.15 59.22
Int8 41.33 64.07 52.17 59.09
one-shot None 43.04 65.55 54.49 55.84
Float16 43.00 65.54 54.62 55.83
Int8 43.03 65.31 54.41 55.44
EN-PT medical zero-shot None 39.32 64.98 48.19 81.77
Float16 39.30 64.97 48.21 81.81
Int8 39.13 64.89 48.35 81.87
one-shot None 39.79 64.36 48.38 79.04
Float16 39.79 64.36 48.38 79.05
Int8 39.74 64.33 48.42 78.95
EN-SW medical zero-shot None 32.61 58.24 53.29 82.85
Float16 32.87 58.41 53.08 82.86
Int8 32.88 58.36 53.08 83.03
one-shot None 34.06 59.12 51.57 82.37
Float16 34.07 59.23 51.45 82.38
Int8 33.97 59.15 51.41 82.42
generic zero-shot None 35.57 57.55 54.91 83.90
Float16 35.54 57.56 54.93 83.92
Int8 35.48 57.46 55.00 83.85
one-shot None 33.79 55.93 56.34 83.23
Float16 33.86 55.96 56.32 83.24
Int8 33.73 55.90 56.22 83.27
SW-EN medical zero-shot None 44.58 64.73 42.71 84.11
Float16 44.48 64.69 42.71 84.17
Int8 44.28 64.60 43.03 84.11
one-shot None 43.66 64.03 42.85 83.92
Float16 43.66 64.01 42.86 83.89
Int8 43.35 63.72 43.23 83.78
generic zero-shot None 42.87 61.59 48.69 82.00
Float16 42.89 61.56 48.63 82.01
Int8 42.83 61.53 48.59 82.00
one-shot None 41.38 59.99 50.85 80.92
Float16 41.40 60.01 50.89 80.95
Int8 41.35 60.01 50.72 80.97
Table 15: Comparison of translation performance of NLLB-200 3.3, across different quantization methods, using CTranslate2. The results show comparable performance.
Lang Context Quantization BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR zero-shot None 39.74 62.51 53.36 66.75
AWQ 39.02 61.97 54.69 64.97
one-shot None 51.93 70.57 43.03 76.76
AWQ 51.83 70.43 43.00 76.19
EN-PT zero-shot None 43.86 67.35 43.95 84.05
AWQ 43.12 66.59 45.58 82.12
one-shot None 45.95 68.82 41.68 86.58
AWQ 45.59 68.57 42.12 86.10
Table 16: Comparison of translation performance of Llama-3 70B, with and without AWQ quantization, using vLLM.